# Nonasymptotic Convergence Rates for Plug-and-Play Methods With MMSE Denoisers

### v3

@@ I-B Plug-and-Play Methods
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Section III-A, we prove that $x\mapsto\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x\|^{2}$ is convex, hence $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.6. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, PGD reaches an $\mathcal{O}(1/\sqrt{k})$-stationary point after $k$ iterations.
@@ II Preliminaries
Many similar versions of this identity are well known in the literature. To our knowledge, this exact formulation has not been explicitly stated.
@@
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.222i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE estimator under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
@@ III Main Results
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient may not exist and is generalized to the subgradient [rockafellar_variational, Chapter 8].
Definition III.4 (Subgradient).
For a function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$, the subgradient $\partial F(x)$ is given by
$\partial F(x)\coloneqq\mathopen{}\mathclose{{\left\{v\in\mathbb{R}^{n}\;\middle|\;\begin{aligned} &F(y)\geq F(x)+\langle v,y-x\rangle,\\ &\text{for all }y\in\mathbb{R}^{n}\end{aligned}}}\right\}$ (39)
Recall that $x^{\star}$ is a stationary point for $F$ if and only if $0\in\partial F(x^{\star})$. This motivates measuring progress by approximate stationarity. In the nonsmooth, non-convex setting, the natural goal is to find an $\varepsilon$-stationary point, i.e. an $x$ satisfying
$\operatorname{dist}(0,\partial F(x))\leq\varepsilon,$ (40)
@@
$\min_{1\leq j\leq k}\operatorname{dist}(0,\partial F(x_{j}))\leq\varepsilon.$ (42)
For our investigation, we prove the following intermediary result, which is a simple improvement over [bohm_variable, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser.
@@
$\min_{2\leq j\leq k+1}\operatorname{dist}(0,\partial F(x_{j}))\leq k^{-1/2}C\sqrt{F(x_{1})-F^{*}},$ (43)
@@
for any $s\in\partial\widetilde{g}(x)$. For iteration $x_{k+1}$, let $s_{k+1}\in\partial g(x_{k+1})$, so that $s_{k+1}+\rho x_{k+1}\in\partial\widetilde{g}(x_{k+1})$. We then have
@@
Setting $s_{k+1}\coloneq\tfrac{1}{\gamma}(x_{k}-x_{k+1})-\nabla f(x_{k})\in\partial g(x_{k+1})$ and adding (III-B) and (III-B),
@@
Next, let $w_{k+1}\coloneqq\tfrac{1}{\gamma}(x_{k+1}-x_{k})+\nabla f(x_{k+1})-\nabla f(x_{k})$, an element of $\partial(f+g)(x_{k+1})$,
@@
using the definition in (44). Taking a square root and recalling that $\operatorname{dist}(0,\partial F(x_{j+1}))\leq\|w_{j+1}\|$ gives us the result. ∎
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [xu_provable] showed that PnP-MMSE converges for any $\gamma\leq 1/L$, but the objective it converges to is not the objective of interest, i.e.,
[…]
but instead it converges to a scaled objective of the form
$F(x)=f(x)+\tfrac{1}{\gamma}\phi_{\mathrm{MMSE}}(x).$ (54)
Furthermore, their convergence is only asymptotic. On the other hand, in this paper, we establish a nonasymptotic convergence result for the objective (53). Our result is summarized in the following theorem.
Theorem III.6.
@@ IV Experiments
Outside the support of the training data, the learned convex potential becomes approximately affine, and its gradient (the denoiser) saturates. To solve for $\phi_{\mathrm{MMSE}}(x)$, we must invert the learned denoiser numerically. This inversion is impossible where the denioser saturates, so we get large values for $\psi_{\theta}^{-1}$ outside the central region. We solve for $\phi_{\mathrm{MMSE}}(x)$ using  [gribonval_characterization, Eq. (8)], so the quadratic $g_{\theta}(\psi_{\theta}^{-1}(x))$ dominates, causing the resulting estimate of the regularizer to curl downward to -$\infty$. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [fang_whats, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
@@
In Figure 5, we plot $\operatorname{dist}(0,{\partial}F(x_{k}))$ over $50$ iterations for the first test image in Figure 4, showing clear sublinear fixed-point convergence, as predicted by Theorem III.6.
@@
In Fig. 8, we plot the stationary residual (41) for 160 iterations and observe behavior consistent with our findings in Theorem III.6
@@ 4 $3\times 10^{-5}$ 5,000
We note that while the restriction of $L<1$ in Theorem III.6 may seem potentially restrictive, any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment, the forward operator, a discretized Radon transform is $L$-smooth with $L\approx 450$. We simply scale the data fidelity term appropriately and still obtain strong empirical results with the desired convergence rate.
@@ V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.6. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, does not transfer to other noise models, so we suspect this flavor of result is unique to the Gaussian noise setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work toward more predictable PnP solvers.
@@ Acknowledgment
The authors would like to thank Stanislas Ducotterd for helpful feedback on this work.

### v4

@@ I-B Plug-and-Play Methods
  • Nonasymptotic convergence of MMSE-PGD via weak convexity. In Section III-A, we prove that $x\mapsto\phi_{\mathrm{MMSE}}(x)+\tfrac{1}{2}\|x\|^{2}$ is convex, hence $\phi_{\mathrm{MMSE}}$ is $1$-weakly convex. With this property we establish (to the best of our knowledge) the first nonasymptotic convergence guarantees for PGD-PnP with an MMSE denoiser in Theorem III.7. In particular, we show that, under an $L$-smooth data-fidelity term $f$ with $L<1$, PGD reaches an $\mathcal{O}(1/\sqrt{k})$-stationary point after $k$ iterations.
@@ II Preliminaries
Many similar versions of this identity are well known in the literature. [^2] Results of this type are often formulated in terms of prox-regularity. We avoid this terminology here, as we don’t use it anywhere else and the proof of the stated result is quite simple. For more detail see [rockafellar_variational, Theorem 13.37].
@@
Let $X\sim p_{X}$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ be independent random variables in $\mathbb{R}^{n}$, and let $Z=X+\epsilon$. Assume $p_{X}$ is non-degenerate.333i.e., there is no pair $v\in\mathbb{R}^{n}$, $c\in\mathbb{R}$ such that $\langle X,v\rangle=c$, almost surely. The MMSE estimator under (2) given by $\psi_{\sigma}(z)\coloneqq\mathbb{E}[X|Z=z]$ has the following properties:
@@ III Main Results
When the objective $F$ is smooth, the map $x\mapsto\nabla F(x)$ is continuous. Therefore, measuring convergence via the gradient norm $\|\nabla F(x_{k})\|$ is natural; small gradient norms certify proximity to a stationary point. In the non-smooth, non-convex setting, the gradient and subgradient (see [rockafellar_variational, Chapter 8]) may not exist and are generalized to the Fréchet subgradient.
Definition III.4 (Fréchet Subgradient).
For a function $F:\mathbb{R}^{n}\to\mathbb{R}\cup\{+\infty\}$ and a point $x$ where $F$ is finite, the Fréchet Subgradient $\widehat{\partial}F(x)$ is given by
$\widehat{\partial}F(x)\coloneqq\mathopen{}\mathclose{{\left\{v\in\mathbb{R}^{n}\;\middle|\;\liminf_{y\to x}\frac{f(x)-f(y)-\langle v,x-y\rangle}{\|x-y\|}}}\right\}.$ (39)
Informally, $\widehat{\partial}F(x)$ is the set of all planes that locally support $F$ at $x$. $x^{\star}$ is a stationary point for $F$ if and only if $0\in\widehat{\partial}F(x^{\star})$, and when $F$ is differentiable, $\widehat{\partial}F(x)=\{\nabla F(x)\}$. This motivates measuring progress by approximate stationarity. In the nonsmooth, non-convex setting, the natural goal is to find an $\varepsilon$-stationary point, i.e. an $x$ satisfying
$\operatorname{dist}(0,\widehat{\partial}F(x))\leq\varepsilon,$ (40)
@@
$\min_{1\leq j\leq k}\operatorname{dist}(0,\widehat{\partial}F(x_{j}))\leq\varepsilon.$ (42)
For our investigation, we prove the following intermediary result, which is a simple improvement over [bohm_variable, Theorem 5.1]. This result will be used to prove convergence of PGD using an MMSE denoiser. For our immediate application, $\widehat{\partial}F(x)$ reduces to $\{\nabla F(x)\}$, but we state the result in the most general form because it provides a unified notion of stationarity that remains valid in nonsmooth and nonconvex settings.
@@
$\min_{2\leq j\leq k+1}\operatorname{dist}(0,\widehat{\partial}F(x_{j}))\leq k^{-1/2}C\sqrt{F(x_{1})-F^{*}},$ (43)
@@
for any $s\in\widehat{\partial}\widetilde{g}(x)$. For iteration $x_{k+1}$, let $s_{k+1}\in\widehat{\partial}g(x_{k+1})$, so that $s_{k+1}+\rho x_{k+1}\in\widehat{\partial}\widetilde{g}(x_{k+1})$. We then have
@@
Setting $s_{k+1}\coloneq\tfrac{1}{\gamma}(x_{k}-x_{k+1})-\nabla f(x_{k})\in\widehat{\partial}g(x_{k+1})$ and adding (III-B) and (III-B),
@@
Next, let $w_{k+1}\coloneqq\tfrac{1}{\gamma}(x_{k+1}-x_{k})+\nabla f(x_{k+1})-\nabla f(x_{k})$, an element of $\widehat{\partial}(f+g)(x_{k+1})$,
@@
using the definition in (44). Taking a square root and recalling that $\operatorname{dist}(0,\widehat{\partial}F(x_{j+1}))\leq\|w_{j+1}\|$ gives us the result. ∎
Remark III.6.
Convergence of PGD using an MMSE denoiser has been previously established for a range of step sizes $\gamma>0$. In particular, [xu_provable] showed that PnP-MMSE converges for any $\gamma\leq 1/L$. Though the MMSE estimator can be written as $\operatorname{prox}_{\phi{\mathrm{MMSE}}}$, it is impossible to evaluate $\operatorname{prox}_{\gamma\phi{\mathrm{MMSE}}}$ for $\gamma\neq 1$. The denoiser fixes the effective step size. Consequently, any change in the step size is really just a rescaling of the data-fidelity term. For this reason, we assume an implicit step size of $1$ and require the Lipschitz constant of the data-fidelity term to be sufficiently small.
Theorem III.7.
@@ IV Experiments
Outside the support of the training data, the learned convex potential becomes approximately affine, and its gradient (the denoiser) saturates. To solve for $\phi_{\mathrm{MMSE}}(x)$, we must invert the learned denoiser numerically. This inversion is impossible where the denoiser saturates, so we get large values for $\psi_{\theta}^{-1}$ outside the central region. We solve for $\phi_{\mathrm{MMSE}}(x)$ using  [gribonval_characterization, Eq. (8)], so the quadratic $g_{\theta}(\psi_{\theta}^{-1}(x))$ dominates, causing the resulting estimate of the regularizer to curl downward to -$\infty$. We emphasize that this behavior arises as a result of the ReLU-approximating architecture, not from the MMSE loss or our implementation. To illustrate this, in Figure 3 we reproduce the setup of [fang_whats, Fig. 2] and simply extend the axis range. The extended view makes the affine-tail/saturation effect apparent.
@@
In Figure 5, we plot $\operatorname{dist}(0,\widehat{\partial}F(x_{k}))$ over $50$ iterations for the first test image in Figure 4, showing clear sublinear fixed-point convergence, as predicted by Theorem III.7.
@@
In Fig. 8, we plot the stationary residual (41) for 160 iterations and observe behavior consistent with our findings in Theorem III.7
@@ 4 $3\times 10^{-5}$ 5,000
We note that while the restriction of $L<1$ in Theorem III.7 may seem potentially restrictive, any $L$-smooth data-fidelity term can be scaled appropriately to meet this condition. For example, in our computed tomography experiment, the forward operator, a discretized Radon transform is $L$-smooth with $L\approx 450$. We simply scale the data fidelity term appropriately and still obtain strong empirical results with the desired convergence rate.
@@ V Conclusion
We presented a novel representation of the MMSE denoiser’s implicit regularizer in Theorem III.2 and used it to derive nonasymptotic convergence results for PnP-PGD using an MMSE denoiser in Theorem III.7. Our analysis is strictly tied to the Gaussian-noise MMSE setting. Tweedie’s formula, for example, does not transfer to other noise models, so we suspect this flavor of result is unique to the Gaussian noise setting. However, we expect analogous nonasymptotic guarantees to be attainable for other PnP schemes that use MMSE denoisers under Gaussian noise, such as PnP-ADMM or the proximal point method. Exploration of these extensions constitutes future work toward more predictable PnP solvers.
@@ Acknowledgment
The authors would like to thank Stanislas Ducotterd and Guillaume Lauga for helpful feedback on this work.